# Interpretable global minima of deep ReLU neural networks on sequentially separable data

### v2

@@ Appendix A Geometry
$\angle(u,v):=\arccos\frac{\langle u,v\rangle}{\left\lVert{u}\right\rVert\left\lVert{v}\right\rVert}.$ (A.1)
@@
$\frac{\langle u,v\rangle}{\left\lVert{u}\right\rVert\left\lVert{v}\right\rVert}\geq\cos\theta.$ (A.3)
@@
$\langle u_{n},v_{j}\rangle=(n-1)\frac{1}{n-1}=1,$ (A.5)
@@
$\angle(u_{n},v_{j})=\arccos\frac{\sqrt{n-1}}{\sqrt{n}},$ (A.6)
giving the expression (3.19) for $\theta_{n}$.
In the proof of Proposition 4.1, we make the assumption $\theta_{*}\geq\frac{\pi}{2}$. Note that a cone based at the origin with axis some unit vector $f\in\mathbb{R}^{n}$ will contain the ball $B_{\delta}(2\delta f)$ if it contains
$\delta f+\delta f^{\perp}$ (A.7)
@@
$=\frac{\left\lVert{\delta f}\right\rVert}{\sqrt{\left\lVert{\delta f}\right\rVert^{2}+\left\lVert{\delta f^{\perp}}\right\rVert^{2}}}$ (A.8)

### v3

@@ Appendix A Geometry
$\angle(u,v):=\arccos\frac{\langle u,v\rangle}{\left\lVert{u}\right\rVert\left\lVert{v}\right\rVert}.$
@@
$\frac{\langle u,v\rangle}{\left\lVert{u}\right\rVert\left\lVert{v}\right\rVert}\geq\cos\theta.$
@@
$\langle u_{n},v_{j}\rangle=(n-1)\frac{1}{n-1}=1,$
@@
$\angle(u_{n},v_{j})=\arccos\frac{\sqrt{n-1}}{\sqrt{n}},$
giving the expression (13) for $\theta_{n}$.
In the proof of Proposition 3, we make the assumption $\theta_{*}\geq\frac{\pi}{2}$. Note that a cone based at the origin with axis some unit vector $f\in\mathbb{R}^{n}$ will contain the ball $B_{\delta}(2\delta f)$ if it contains
$\delta f+\delta f^{\perp}$
@@
$=\frac{\left\lVert{\delta f}\right\rVert}{\sqrt{\left\lVert{\delta f}\right\rVert^{2}+\left\lVert{\delta f^{\perp}}\right\rVert^{2}}}$